n a jump spike, Mark Angel slams the volleyball from a height of 2.3 m with a speed of 18 m/s at 24° below the horizontal. How far on the opposite side will the ball land?


1
Expert's answer
2022-03-24T16:36:59-0400

Answer

"d=\n\\frac{\n18cos24}{9.8}\n\u200b\t\n(18sin24+\\sqrt{ ( \n(18sin24) \n2\n +2(9.8(2.3)\n\u200b\t\n} )\n=29 m"